### v1

# Optimal state estimation: Turnpike analysis and performance results

Abstract
In this paper, we introduce turnpike arguments in the context of optimal state estimation. In particular, we show that the optimal solution of the state estimation problem involving all available past data serves as turnpike for the solutions of truncated problems involving only a subset of the data. We consider two different mathematical characterizations of this phenomenon and provide corresponding sufficient conditions that rely on strict dissipativity and decaying sensitivity. As second contribution, we show how a specific turnpike property can be used to establish performance guarantees when approximating the optimal solution of the full problem by a sequence of truncated problems, and we show that the resulting performance (both averaged and non-averaged) is approximately optimal with error terms that can be made arbitrarily small by an appropriate choice of the horizon length. In addition, we discuss interesting implications of these results for the practically relevant case of moving horizon estimation and illustrate our results with a numerical example.
I Introduction
Reconstructing the internal state trajectory of a dynamical system based on a batch of measured input-output data is an important problem of high practical relevance. This can be accomplished, for example, by solving an optimization problem to find the best state and disturbance trajectories that minimize a suitably defined cost function depending on the measurement data. If all available data is taken into account, this corresponds to the full information estimation (FIE) problem. However, if the data set or the underlying model is very large or only a limited amount of computation time or resources is available (as is the case with, e.g., online state estimation), the optimal solution to the FIE problem is usually difficult (or even impossible) to compute in practice. For this reason, it is essential to find a reasonable approximation, which can be done, e.g., by means of a sequence of truncated optimal estimation problems, each of which uses only a limited time window of the full data set. In the case of online state estimation, this corresponds to moving horizon estimation (MHE), where in each discrete time step an optimal estimation problem is solved with a data set of fixed size.
Current research in the field of MHE is primarily concerned with stability and robustness guarantees, see, e.g., [1, Ch. 4] and [2, 3, 4, 5]. These works essentially show that under suitable detectability conditions, the estimation error of MHE converges to a neighborhood of the origin, whose size depends on the true unknown disturbance. However, results on the actual performance of MHE, and in particular on the approximation accuracy and performance loss compared to the desired optimal (but unknown) FIE solution, are lacking.
Whereas performance guarantees for state estimators are generally rather rare and usually restricted to linear systems (cf., e.g., [6, 7]), they often play an important role in nonlinear optimal control, especially when the overall goal is an economic one. Corresponding results usually employ a turnpike property of the underlying optimal control problem, cf. [8, 9]. This property essentially implies that optimal trajectories stay close to an optimal equilibrium (or in general an optimal time-varying reference) most of the time, which is regarded as the turnpike. Turnpike-related arguments are an important tool for assessing the closed-loop performance of model predictive controllers with general economic costs on finite and infinite horizons, cf. [10, 11, 12]. Necessary and sufficient conditions for the presence of the turnpike phenomenon in optimal control problems are discussed in, e.g., [13, 14] and are usually based on dissipativity, controllability, and suitable optimality conditions.
Our contribution is twofold. First, in Section III we introduce turnpike arguments in the context of optimal state estimation, which is a novelty in itself. In particular, we show that the optimal FIE solution involving all past data serves as turnpike for the solutions of truncated state estimation problems. We consider two different mathematical characterizations of this phenomenon and provide corresponding sufficient conditions: strict dissipativity (Theorem 1) and decaying sensitivity (Theorem 2). Second, in Section IV, we provide novel performance guarantees for optimal state estimation. To this end, we construct a candidate for approximating the FIE solution by a sequence of truncated problems (which can be solved in parallel for offline estimation). Assuming that the optimal state estimation problem exhibits a certain turnpike property, we show that the performance of this candidate is approximately optimal, both averaged and non-averaged, with error terms that can be made arbitrarily small by an appropriate choice of horizon length. In addition, we discuss interesting implications of these results for MHE.
Notation
We denote the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a\in\mathbb{I}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]\subset\mathbb{I}$ by $\mathbb{I}_{[a,b]}$. The Euclidean norm of a vector $x\in\mathbb{R}^{n}$ is denoted by $|x|$, and the weighted norm with respect to a positive definite matrix $Q=Q^{\top}$ by $|x|_{Q}=\sqrt{x^{\top}Qx}$. The maximum eigenvalue of $Q$ is denoted by $\lambda_{\max}(Q)$. We refer to a sequence $\{x_{j}\}_{j=a}^{b}$, $x_{j}\in\mathbb{R}^{n}$, $j\in\mathbb{I}_{[a,b]}$ with $x_{a:b}$. Finally, we recall that a function $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is of class $\mathcal{K}$ if it is continuous, strictly increasing, and satisfies $\alpha(0)=0$; if additionally $\alpha(r)=\infty$ for $r\rightarrow\infty$, it is of class $\mathcal{K}_{\infty}$. By $\mathcal{L}$, we refer to the class of functions $\theta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ that are continuous, non-increasing, and satisfy $\lim_{s\rightarrow\infty}\theta(s)=0$, and by $\mathcal{KL}$ to the class of functions $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ with $\beta(\cdot,s)\in\mathcal{K}$ and $\beta(r,\cdot)\in\mathcal{L}$ for any fixed $s\in\mathbb{R}_{\geq 0}$ and $r\in\mathbb{R}_{\geq 0}$, respectively.
II Problem Setup
II-A System description
We consider the following system
${x}_{t+1}$ $=f(x_{t},u_{t},w_{t}),$ (1a)
$y_{t}$ $=h(x_{t},u_{t})+v_{t}$ (1b)
with discrete time $t\in\mathbb{I}_{\geq 0}$, state $x_{t}\in\mathbb{R}^{n}$, (known) control input $u_{t}\in\mathbb{R}^{m}$, (unknown) disturbances and noise $w_{t}\in\mathbb{R}^{q}$ and $v_{t}\in\mathbb{R}^{p}$, and noisy output measurement $y_{t}\in\mathbb{R}^{p}$. The functions $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{q}\rightarrow\mathbb{R}^{n}$ and $h:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{p}$ define the system dynamics and output equation, which we assume to be continuous.
In the following, we consider the general case where we may know some additional information in the form of $(x_{t},u_{t},w_{t},v_{t})\in\mathcal{X}\times\mathcal{U}\times\mathcal{W}\times\mathcal{V}=:\mathcal{C},\ t\in\mathbb{I}_{\geq 0}$ with known sets $\mathcal{X}\subseteq\mathbb{R}^{n}$, $\mathcal{U}\subseteq\mathbb{R}^{m}$, $\mathcal{W}\subseteq\mathbb{R}^{q}$, $\mathcal{V}\subseteq\mathbb{R}^{p}$, and furthermore, that $(x_{t},u_{t},w_{t},v_{t})\in\mathcal{C}\Rightarrow x_{t+1}\in\mathcal{X}$ for all $t\in\mathbb{I}_{\geq 0}$. Such knowledge typically arises from the physical nature of the system (e.g., non-negativity of certain physical quantities), the incorporation of which can significantly improve the estimation results, cf. [1, Sec. 4.4] and compare also [4].
II-B Optimal state estimation
For ease of notation, we define the input-output data tuple $d_{t}:=(u_{t},y_{t})$ obtained from the system (1) at time $t\in\mathbb{I}_{\geq 0}$. Now, consider a given batch of input-output data $d_{0:T}$ of length $T+1$ for some $\mathbb{I}_{\geq 0}$. We aim to compute the state and disturbance sequences $\hat{x}_{0:T}$ and $\hat{w}_{0:T-1}$ that are optimal in the sense that they minimize a cost function involving the full data set $d_{0:T}$. In particular, we consider the following (FIE) optimization problem $P_{T}(d_{0:T})$:
$\min_{\text{\footnotesize$\begin{matrix}\hat{x}_{0:T}\\ \hat{w}_{0:T-1}\end{matrix}$}}\ J_{T}(\hat{x}_{0:T},\hat{w}_{0:T-1};d_{0:T})$ (2a)
s.t. $\hat{x}_{j+1}=f(\hat{x}_{j},u_{j},\hat{w}_{j}),\ j\in\mathbb{I}_{[0,T-1]},$ (2b)
$\hat{x}_{j}\in\mathcal{X},\ y_{j}-h(\hat{x}_{j},u_{j})\in\mathcal{V},\ j\in\mathbb{I}_{[0,T]},$ (2c)
$\hat{w}_{j}\in\mathcal{W},\ j\in\mathbb{I}_{[0,T-1]}.$ (2d)
For convenience, we define the combined sequence $\hat{z}_{0:T}$ satisfying $\hat{z}_{j}=(\hat{x}_{j},\hat{w}_{j})$ for $j\in\mathbb{I}_{[0,T-1]}$ and $\hat{z}_{T}=(\hat{x}_{T},0)$. The constraints (2b)–(2d) enforce the prior knowledge about the system model, the domain of the true trajectories, and the disturbances/noise (note that feasibility is always guaranteed due to our standing assumptions). In (2a), we consider the cost function
$J_{T}(\hat{x}_{0:T},\hat{w}_{0:T-1};d_{0:T})=\sum_{j=0}^{T-1}l(\hat{x}_{j},\hat{w}_{j};d_{j})+g(\hat{x}_{T};d_{T})$ (3)
with continuous stage cost $l:\mathcal{X}\times\mathcal{W}\times\mathcal{U}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{\geq 0}$ and terminal cost $g:\mathcal{X}\times\mathcal{U}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{\geq 0}$. Note that this is a generalization of classical designs for state estimation, where $l$ and $g$ are positive definite in the disturbance input $\hat{w}$ and the fitting error $y-h(\hat{x},u)$ (cf., e.g., [1, Ch. 4]), and it particularly includes the practically relevant case of quadratic stage and terminal cost
$l(x,w;d)=|w|^{2}_{Q}+|y-h(x,u)|^{2}_{R}$ (4)
and
$g(x;d)=|y-h(x,u)|^{2}_{G},$ (5)
respectively, where $Q,R,G$ are positive definite weighting matrices. However, our results presented in Section III also hold for more general cost functions $l$ and $g$, which allow the objective (3) to be tailored to the specific problem at hand.
In the state estimation context, a cost function with terminal cost as in (5) is usually referred to as the filtering form of the state estimation problem. To simplify the analysis and notation, the most recent works on FIE/MHE theory often consider a cost function with $g=0$ (i.e., without terminal cost), which corresponds to the prediction form of the state estimation problem, cf., e.g., [4, 3, 2, 5], and see also [1, Ch. 4] for a discussion on this topic.
The FIE problem $P_{T}$ in (2) is a parametric nonlinear program, the solution of which solely depends on the (input-output) data provided, i.e., the sequence $d_{0:T}$. We characterize solutions to $P_{T}$ using a generic mapping $\zeta_{T}$ that maps data to optimal states and disturbance inputs:
$z^{*}_{j}=\zeta_{T}(j,d_{0:T}),\ j\in\mathbb{I}_{[0,T]}.$ (6)
This yields the value function $V_{T}(d_{0:T})=J_{T}(z^{*}_{0:T};d_{0:T})$. However, if the data set (in particular, $T$) or the underlying model is very large or the computations are limited in terms of time or resources (as is the case with online state estimation, for example), solving the FIE problem $P_{T}$ for the optimal solution is usually difficult (or even impossible) in practice. For this reason, we also consider solutions of the truncated optimal state estimation problem ${P_{N}(d_{\tau:\tau+N})}$ using a fixed horizon length $N\in\mathbb{I}_{[0,T]}$, i.e., the problem $P_{T}$ in (2) with $T$ replaced by $N$ and the truncated data sequence $d_{\tau:\tau+N}$ for some $\tau\in\mathbb{I}_{[0:T-N]}$. We characterize corresponding solutions using the generic solution mapping $\zeta_{N}$:
$\hat{z}_{\tau+j}^{*}:=\zeta_{N}(j,d_{\tau:\tau+N}),\ j\in\mathbb{I}_{[0,N]},$ (7)
with the value function $V_{N}(d_{\tau:\tau+N}){\,=\,}J_{N}(\hat{z}_{\tau:\tau+N}^{*};d_{\tau:\tau+N})$. In the following, we will investigate how the solution $\hat{z}^{*}_{\tau:\tau+N}$ (7) of the truncated estimation problem $P_{N}(d_{\tau:\tau+N})$ behaves compared to the optimal FIE solution $z^{*}_{0:T}$ (6) on the interval $\mathbb{I}_{[\tau:\tau+N]}$ (Section III), how it can be suitably approximated on $\mathbb{I}_{[0,T]}$ using a sequence of truncated problems $P_{N}$, and how good this approximation actually is (Section IV).
III Turnpike in optimal state estimation
Optimal state estimation problems (such as $P_{T}$ in (2)) can generally be interpreted as optimal control problems using $\hat{w}$ as the control input (compare also [1, Sec. 4.2.3] and [15, Sec. 4]). In particular, a cost function (3) which is positive definite in the estimated disturbance and the fitting error (as, e.g., in (4) and (5)) can be regarded as an economic output tracking cost associated with the ideal reference $(w^{\mathrm{r}}_{j},y^{\mathrm{r}}_{j})=(0,y_{j})$, $j\in\mathbb{I}_{[0,T]}$. This reference, however, is unreachable (assuming that the system is detectable), as it is generally impossible to attain zero cost $V_{T}(d_{0:T})=0$ (except for the special case where $y_{0:T}$ corresponds to an output sequence of (1) under zero disturbances $w_{0:T-1}\equiv 0$, $v_{0:T}\equiv 0$). For unreachable references, on the other hand, it is known that the corresponding optimal control problem exhibits the turnpike property with respect to the best reachable reference [16], which suggests that a similar phenomenon can also be expected in optimal state estimation problems. In the following, we provide a simple example that supports this intuition (Section III-A). Then, we consider two different turnpike characterizations and provide corresponding sufficient conditions that rely on strict dissipativity (Section III-B) and decaying sensitivity (Section III-C).
III-A Motivating example
Suppose that data $y_{0:T}$ is measured from the system $x_{t+1}=x_{t}+w_{t}$, $x_{0}=1$, $y_{t}=x_{t}+v_{t}$, where $w_{t}=v_{t}=1$ for all $t\in\mathbb{I}_{\geq 0}$. We consider the optimal estimation problem $P_{T}(d_{0:T})$ with quadratic stage and terminal cost (4) and (5) using $Q=R=G=1$, $T=70$, and solutions of the truncated problem $P_{N}(d_{\tau:\tau+N})$ for different values of $N$ and $\tau$.
Figure 1: Difference $|\hat{x}_{j}^{*}-{x}_{j}^{*}|$, $j\in\mathbb{I}_{[\tau,\tau+N]}$ for $N=5$ (cyan), $N=10$ (magenta), $N=15$ (blue), and $N=20$ (red) for different values of $\tau$.
Figure 1 shows the difference between the optimal state $x^{*}_{j}$ and the solution of the truncated problem $\hat{x}^{*}_{j}$ for $j\in\mathbb{I}_{[\tau,\tau+N]}$. One can clearly see that the optimal reference $x^{*}_{0:T}$ serves as turnpike for the solution $\hat{x}^{*}_{\tau:\tau+N}$. In particular, $\hat{x}^{*}_{\tau:\tau+N}$ is constructed from three pieces: an initial transient where $\hat{x}^{*}$ converges to the turnpike ${x}^{*}$, a large phase where $\hat{x}^{*}$ stays near the turnpike ${x}^{*}$, and a transient at the end of the horizon where $\hat{x}^{*}$ diverges from the turnpike ${x}^{*}$. Figure 1 indicates that these transients111 In the turnpike-related literature, the left transient is usually referred to as approaching arc and the right transient as leaving arc. are independent of the horizon length $N$. For the special cases at the boundaries ($\tau=0$ and $\tau=T-N$), they only occur on one side (for an explanation, see Remark 2 below). Note that a similar behavior is also observed (but not shown due to lack of space) for the input difference $\hat{w}^{*}_{j}-{w}^{*}_{j}$, $j\in\mathbb{I}_{[\tau,\tau+N]}$.
III-B Turnpike under strict dissipativity
In this subsection, we consider the terminal cost $g(x;d)=0$ to simplify the analysis.
Assumption 1 (Strict dissipativity)
Consider the data sequence $d_{0:T}$ and the optimal reference solution $z_{j}^{*}=\zeta_{T}(j,d_{0:T})$. The optimal state estimation problem (2) is strictly dissipative with respect to the supply rate
$s_{j}(x,w)=l(x,w;d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j}),$ (8)
i.e., there exists a time-varying storage function $\lambda_{j}:\mathbb{I}_{[0,T]}\times\mathcal{X}\rightarrow\mathbb{R}$ and $\alpha\in\mathcal{K}_{\infty}$ such that
$\lambda_{j+1}(f(x,u_{j},w))-\lambda_{j}(x_{j})\leq-\alpha(|z-z_{j}^{*}|)+s_{j}(x,w)$ (9)
for all $j\in\mathbb{I}_{[0,T-1]}$ and all $z=(x,w)\in\mathcal{X}\times\mathcal{W}$. Moreover, there exists $\alpha_{\lambda}\in\mathcal{K}_{\infty}$ such that $|\lambda_{j}(x)|\leq\alpha_{\lambda}(|x-x_{j}^{*}|)$ uniformly for all $j\in\mathbb{I}_{[0,T]}$.
Strict dissipativity is a standard assumption in the context of (economic) model predictive control, usually invoked to establish turnpike behavior of open loop optimal control problems and obtain stability and performance guarantees of the resulting closed loop, see, e.g., [17, 12, 16]. The following result shows that under strict dissipativity, the solutions of the truncated problems $P_{N}$ are close to the optimal solution of the FIE problem $P_{T}$ for most of the time, which hence constitutes the turnpike associated with the optimal estimation problem.
Theorem 1 (Strict dissipativity implies turnpike)
Let $\mathcal{X}$ be compact and suppose that Assumption 1 is satisfied. Consider the optimal FIE solution ${z}^{*}_{j}=\zeta_{T}(j,d_{0:T})$, $j\in\mathbb{I}_{[0,T]}$ with some given $d_{0:T}$ and $T\in\mathbb{I}_{\geq 0}$. Then, there exists $\sigma\in\mathcal{L}$ such that for all $P\in\mathbb{I}_{[0,N]}$, at least $P$ points $j\in\mathbb{I}_{[0,N-1]}$ corresponding to the solution $\hat{z}^{*}_{\tau+j}=\zeta_{N}(j,d_{\tau:\tau+N})$ satisfy
$|\hat{z}^{*}_{\tau+j}-z^{*}_{\tau+j}|\leq\sigma(N-P)$ (10)
for all $N\in\mathbb{I}_{[0,T]}$, $\tau\in\mathbb{I}_{[0,T-N]}$, and all possible data $d_{0:T}$.
Proof:
We consider the rotated stage cost
$L_{j}(x,w)=$ $\ l(x,w;d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j})$
$\ +\lambda_{j}(x)-\lambda_{j+1}(f(x,u_{j},w)),\ j\in\mathbb{I}_{[0,T-1]}$
for any point $z=(x,w)\in\mathcal{X}\times\mathcal{W}$. From strict dissipativity (Assumption 1), we know that
$L_{j}(x,w)\geq\alpha(|z-z_{j}^{*}|),\ j\in\mathbb{I}_{[0,T-1]},$ (11)
which implies that the rotated stage cost is positive definite with respect to the difference $|z-z_{j}^{*}|$ for all $j\in\mathbb{I}_{[0:T-1]}$ and any $z\in\mathcal{X}\times\mathcal{W}$.
Now consider some $N\in\mathbb{I}_{[0,T]}$, $\tau\in\mathbb{I}_{[0,T-N]}$, and the solution $\hat{z}^{*}_{\tau+j}=\zeta_{N}(j,d_{\tau:\tau+N})$, $j\in\mathbb{I}_{[0,N]}$. Evaluating the rotated cost function along $\hat{z}^{*}_{\tau:\tau+N}$ yields
$\sum_{j=\tau}^{\tau+N-1}L_{j}(\hat{x}^{*}_{j},\hat{w}^{*}_{j})=$ $\ \sum_{j=\tau}^{\tau+N-1}l(\hat{x}^{*}_{j},\hat{w}^{*}_{j};d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j})$
$\ +\lambda_{\tau}(\hat{x}^{*}_{\tau})-\lambda_{\tau+N}(\hat{x}^{*}_{\tau+N}).$
The application of (11) then leads to
$\sum_{j=\tau}^{\tau+N-1}l(\hat{x}^{*}_{j},\hat{w}^{*}_{j};d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j})$
$\geq\lambda_{\tau+N}(\hat{x}^{*}_{\tau+N})-\lambda_{\tau}(\hat{x}^{*}_{\tau})+\sum_{j=\tau}^{\tau+N-1}\alpha(|\hat{z}^{*}_{j}-z^{*}_{j}|).$ (12)
Due to the fact that $\mathcal{X}$ is compact and the storage function is bounded (Assumption 1), we can define $C_{\lambda}:=\max\limits_{x_{1},x_{2}\in\mathcal{X}}\alpha_{\lambda}(|x_{1}-x_{2}|)$ such that $|\lambda_{j}(x)|\leq C_{\lambda}$ for all $j\in\mathbb{I}_{[\tau,\tau+N]}$ and all $x\in\mathcal{X}$. We claim that the turnpike property holds with
$\sigma(r)=\alpha^{-1}\left(\frac{2C_{\lambda}}{r}\right).$ (13)
For the sake of contradiction, suppose not. Then, at least $N-P+1$ points $j\in\mathbb{I}_{[0,N-1]}$ satisfy $|\hat{z}^{*}_{\tau+j}-z^{*}_{\tau+j}|>\sigma(N-P)$. Hence, from (12) it follows that
$\sum_{j=\tau}^{\tau+N-1}l(\hat{x}^{*}_{j},\hat{w}^{*}_{j};d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j})$
$\quad>-2C_{\lambda}+(N-P)\alpha(\sigma(N-P))=0,$ (14)
where the last equality follows from (13). However, the inequality in (14) contradicts optimality of $\hat{z}^{*}_{\tau:\tau+N}$ since $\sum_{j=\tau}^{\tau+N-1}l(\hat{x}^{*}_{j},\hat{w}^{*}_{j};d_{j})-l(x_{j}^{*},w_{j}^{*};d_{j})=V_{N}(d_{\tau:\tau+N})-J_{N}(z^{*}_{\tau:\tau+N};d_{\tau:\tau+N})\leq 0$, which proves our claim. ∎
Remark 1
Theorem 1 and its proof follow the lines of classical turnpike results in the context of optimal control, e.g., [18, Th. 1]. However, because the initial state is free in the estimation problem and not fixed as typical in control, we can avoid the need for an additional reachability/controllability property to bound the value function $V_{N}(d_{\tau:\tau+N-1})$, simply by exploiting the fact that the trajectory $z^{*}_{\tau:\tau+N}$ is a feasible candidate solution.
III-C Turnpike under a decaying sensitivity condition
In the previous section, we established turnpike behavior in optimal state estimation problems using strict dissipativity. The property guaranteed by Theorem 1 is, however, rather weak in the sense that we actually do not know which points of the solution $\hat{z}_{j}$ are close to the turnpike $z^{*}_{j}$ and which not. We can, however, strengthen this result under a decaying sensitivity condition, cf. [19, 20, 21] and see also [22, 23].
To this end, consider the auxiliary nonlinear program $\bar{P}(\bar{d}_{0:N},x^{\mathrm{i}},x^{\mathrm{t}})$ parameterized by data $(\bar{d}_{0:N},x^{\mathrm{i}},x^{\mathrm{t}})$, where $\bar{d}_{j}=(\bar{u}_{j},\bar{y}_{j})\in\mathcal{U}\times\mathbb{R}^{p}$, $j\in\mathbb{I}_{[0,N]}$:
$\min_{\text{\footnotesize$\begin{matrix}\bar{x}_{0:N}\\ \bar{w}_{0:N-1}\end{matrix}$}}\ \sum_{j=0}^{N-1}l(\bar{x}_{j},\bar{w}_{j};\bar{d}_{j})+g(\bar{x}_{N};\bar{d}_{N})$ (15a)
s.t. $\bar{x}_{0}=x^{\mathrm{i}}$ (15b)
$\bar{x}_{N}=x^{\mathrm{t}}$ (15c)
$\bar{x}_{j+1}=f(\bar{x}_{j},\bar{u}_{j},\bar{w}_{j}),\ j\in\mathbb{I}_{[0,N-1]}$ (15d)
$\bar{x}_{j}\in\mathcal{X},\ \bar{y}_{j}-h(\bar{x}_{j},\bar{u}_{j})\in\mathcal{V},\ j\in\mathbb{I}_{[0,N]},$ (15e)
$\bar{w}_{j}\in\mathcal{W},\ j\in\mathbb{I}_{[0,N-1]}.$ (15f)
Let $\bar{z}_{j}:=(\bar{x}_{j},\bar{w}_{j})$, $j\in\mathbb{I}_{[0,N-1]}$, and $\bar{z}_{N}:=(\bar{x}_{N},0)$. We denote solutions to this optimization problem as $\bar{z}^{*}_{j}=\bar{\zeta}_{N}(j,\bar{d}_{0:N},x^{\mathrm{i}},x^{\mathrm{t}})$, $j\in\mathbb{I}_{[0,N]}$.
Lemma 1
Any minimizer of ${P}(d_{0:T})$ is a minimizer of $\bar{P}(d_{\tau:\tau+N},x^{*}_{\tau},x^{*}_{\tau+N})$ for any $N\in\mathbb{I}_{[0,T]}$, $\tau\in\mathbb{I}_{[0,T-N]}$.
Proof:
The proof relies on the principle of optimality. First, note that $\bar{P}(d_{\tau:\tau+N},x^{*}_{\tau},x^{*}_{\tau+N})$ is feasible because the candidate solution $z^{*}_{\tau:\tau+N}$ satisfies the constraints (15b)–(15f) with $\bar{d}_{0:N}=d_{\tau:\tau+N}$. Let $\bar{z}^{*}_{\tau+j}:=\bar{\zeta}_{N}(j,d_{\tau:\tau+N},x^{*}_{\tau},x^{*}_{\tau+N})$. Suppose the claim is false. Then,
$\sum_{j=\tau}^{\tau+N-1}l(z_{j}^{*};d_{j})+g(x^{*}_{\tau+N};d_{\tau+N})$
$>\sum_{j=\tau}^{\tau+N-1}l(\bar{z}^{*}_{j};d_{j})+g(\bar{x}^{*}_{\tau+N};d_{\tau+N}).$
To both sides of this inequality, we add the cost terms
$\sum_{j=0}^{\tau-1}l(z_{j}^{*};d_{j})+\sum_{j=\tau+N}^{T-1}l(z_{j}^{*};d_{j})-g(x^{*}_{\tau+N};d_{\tau+N})+g(x^{*}_{T};d_{T}),$
which yields
$V(d_{0:T})>$ $\sum_{j=0}^{\tau-1}l(z_{j}^{*};d_{j})+\sum_{j=\tau}^{\tau+N-1}l(\bar{z}^{*}_{j};d_{j})+\sum_{j=\tau+N}^{T-1}l(z_{j}^{*};d_{j})$
$+g(\bar{x}^{*}_{\tau+N};d_{\tau+N})+g(x^{*}_{T};d_{T})-g(x^{*}_{\tau+N};d_{\tau+N}).$
Because of the terminal constraint (15c), we have that $\bar{x}^{*}_{\tau+N}=x^{*}_{\tau+N}$, which leads to
$V(d_{0:T})$ $>\sum_{j=0}^{\tau-1}l(z_{j}^{*};d_{j})+\sum_{j=\tau}^{\tau+N-1}l(\bar{z}^{*}_{j};d_{j})+\sum_{j=\tau+N}^{T-1}l(z_{j}^{*};d_{j})$
$\quad+g(x^{*}_{T};d_{T}).$ (16)
Now consider the sequence $\xi_{0:T}$ defined as
$\xi_{j}=(\chi_{j},\omega_{j}):=\begin{cases}{z}^{*}_{j},&j\in\mathbb{I}_{[0,\tau-1]}\\ \bar{z}^{*}_{j},&j\in\mathbb{I}_{[\tau,\tau+N-1]}\\ {z}^{*}_{j},&j\in\mathbb{I}_{[\tau+N,T]}.\end{cases}$ (17)
Then, (16) can be re-written as
$V(d_{0:T})>\sum_{j=0}^{T-1}l(\xi_{j}^{*};d_{j})+g(\chi_{T}^{*};d_{T}).$ (18)
Note that $\xi_{0:T}$ (17) satisfies the constraints in (2) for all $j\in\mathbb{I}_{[0,T]}$ due to the constraints (15b) and (15c), and hence constitutes a valid candidate solution for the optimal estimation problem $P(d_{0:T})$. However, (18) contradicts minimality of $V(d_{0:T})$, which finishes this proof. ∎
We make the following assumption.
Assumption 2 (Decaying sensitivity)
There exists $\beta\in\mathcal{KL}$ such that for any $x^{\mathrm{i}}_{1},x^{\mathrm{i}}_{2}\in\mathcal{X}$ and $x^{\mathrm{t}}_{1},x^{\mathrm{t}}_{2}\in\mathcal{X}$ and any data $d_{\tau:\tau+N}$ for which (15) is feasible, it holds that
$|\bar{\zeta}_{N}(j,d_{\tau:\tau+N},x^{\mathrm{i}}_{1},x^{\mathrm{t}}_{1})-\bar{\zeta}_{N}(j,d_{\tau:\tau+N},x^{\mathrm{i}}_{2},x^{\mathrm{t}}_{2})|$
$\leq\beta(|x^{\mathrm{i}}_{1}-x^{\mathrm{i}}_{2}|,{j})+\beta(|x^{\mathrm{t}}_{1}-x^{\mathrm{t}}_{2}|,{N-j})$ (19)
for all $j\in\mathbb{I}_{[0,N]}$, and all $N\in\mathbb{I}_{[0,T]}$ and $\tau\in\mathbb{I}_{[0,T-N]}$.
Decaying sensitivity is a quite natural property of parametric nonlinear programs that characterizes how much perturbations in the data at one stage influence the solution of the optimization problem at another stage, cf. [19, 20]. In fact, it has been shown that exponentially decaying sensitivity is present under certain observability and controllability conditions, see [21], and compare also [23]. In our case, the property (19) refers to two different data sets that only differ in $x^{\mathrm{i}}_{1},x^{\mathrm{i}}_{2}$ and $x^{\mathrm{t}}_{1},x^{\mathrm{t}}_{2}$, but otherwise contain exactly the same data; consequently, the bound given by the right-hand side in (19) involves only those terms.
Theorem 2 (Decaying sensitivity implies turnpike)
Let $\mathcal{X}$ be compact and suppose that Assumption 2 is satisfied. Consider the optimal FIE solution ${z}^{*}_{j}=\zeta_{T}(j,d_{0:T})$, $j\in\mathbb{I}_{[0,T]}$ with some given $d_{0:T}$ and $T\in\mathbb{I}_{\geq 0}$. Then, there exists $C>0$ such that the solution $\hat{z}^{*}_{\tau+j}=\zeta_{N}(j,d_{\tau:\tau+N})$ satisfies
$|\hat{z}_{\tau+j}^{*}-z_{\tau+j}^{*}|\leq\beta(C,j)+\beta(C,N-j)$ (20)
for all $j\in\mathbb{I}_{[0,N]}$, $\tau\in\mathbb{I}_{[0,T-N]}$, $N\in\mathbb{I}_{[0,T]}$, and $T\in\mathbb{I}_{\geq 0}$, and all possible data $d_{0:T}$.
Proof:
Theorem 2 is a simple consequence of the fact that the solution $\hat{z}_{\tau:\tau+N}^{*}$ is a minimizer of $\bar{P}(d_{\tau:\tau+N},\hat{z}^{*}_{\tau},\hat{z}^{*}_{\tau+N})$, and $z_{\tau:\tau+N}^{*}$ of $\bar{P}(d_{\tau:\tau+N},x^{*}_{\tau},x^{*}_{\tau+N})$ (by Lemma 1), i.e., the same nonlinear program with slightly different data (the initial and terminal condition). The application of (19) in Assumption 2 then leads to $|\hat{z}_{\tau+j}^{*}-z_{\tau+j}^{*}|\leq\beta(|\hat{x}^{*}_{\tau}-x^{*}_{\tau}|,j)+\beta(|\hat{x}^{*}_{\tau+N}-x^{*}_{\tau+N}|,N-j)$ for all $j\in\mathbb{I}_{[0,N]}$ and $\tau\in\mathbb{I}_{[0,T-N]}$. By compactness of $\mathcal{X}$, there exists a constant $C$ such that $|x_{1}-x_{2}|\leq C$ for all $x_{1},x_{2}\in\mathcal{X}$. In combination, we obtain (20), which finishes this proof. ∎
Remark 2 (Approaching and leaving arcs)
Different initial and terminal conditions in the auxiliary problem $\bar{P}$ account for the influence of the neglected data in the truncated problem $P_{N}$ compared to the FIE problem $P_{T}$. For $\tau\in\mathbb{I}_{[1,T-N-1]}$, neglected data involves both the past data $d_{0:\tau-1}$ and the future data $d_{\tau+N+1:T}$, which is the reason why solutions of the truncated problems exhibit both a left approaching arc and a right leaving arc (compare also the motivating example in Section III-A). For $\tau=0$ and $\tau=T-N$, however, the problems $P_{N}$ and $P_{T}$ (in particular, their data) actually half-side coincide, which is why no approaching and leaving arcs appear on the left and right side in Figure 1. These special cases can be theoretically covered by considering modified problems $\bar{P}$ (with either the initial or terminal constraint being enforced) and corresponding decaying sensitivity assumptions, leading to improved bounds; namely, for $\tau=0$, we can derive $|\hat{z}_{j}^{*}-z_{j}^{*}|\leq\beta(C,N-j),\ j\in\mathbb{I}_{[0,N]}$, and for $\tau=T-N$: $|\hat{z}_{T-N+j}^{*}-z_{T-N+j}^{*}|\leq\beta(C,j),\ j\in\mathbb{I}_{[0,N]}$.
IV Performance of optimal state estimation
In this section, we show how the FIE problem $P_{T}$ can be suitably approximated by a sequence of truncated problems $P_{N}$, and that this approximation is in fact approximately optimal with respect to the full problem (Section IV-A). Furthermore, we discuss implications for the practically relevant case of MHE for online state estimation (Section IV-B). The key condition for the results in this section is a turnpike property of certain form.
Assumption 3 (Turnpike behavior)
There exists $\beta\in\mathcal{KL}$ and $C>0$ such that for all $T\in\mathbb{I}_{\geq 0}$ and $N\in\mathbb{I}_{[0,T]}$, the solutions of the problems $P_{T}(d_{0:T})$ and $P_{N}(d_{\tau:\tau+N})$ satisfy
$|\hat{z}_{\tau+j}^{*}-z_{\tau+j}^{*}|\leq\begin{cases}\beta(C,N-j),&\tau=0\\ \beta(C,j)+\beta(C,N-j),&\tau\in\mathbb{I}_{[1,T-N-1]}\\ \beta(C,j),&\tau=T-N\end{cases}$
for all $j\in\mathbb{I}_{[0,N]}$, $\tau\in\mathbb{I}_{[0,T-N]}$, and all possible data $d_{0:T}$.
Note that Assumption 3 is satisfied if the optimal state estimation problem exhibits a decaying sensitivity property, see Theorem 2 and Remark 2.
IV-A Approximation properties and performance guarantees
We aim to approximate the FIE solution $z_{0:T}^{*}$ using a sequence of truncated problems $P_{N}$ of length $N\in\mathbb{I}_{\geq 0}$ (where for simplicity we assume that $N$ is an even number). Since the individual truncated problems are completely decoupled from each other, computing this approximation can be parallelized and hence has the potential to significantly save time and resources. Moreover, we show how a careful concatenation of the solutions to the truncated problems leads to a sequence that is approximately optimal with respect to the (unknown) optimal FIE solution, with error terms that can be made arbitrarily small. This is very interesting from a theoretical point of view, but also practically relevant in, e.g., large data assimilation problems that appear in geophysics and environmental sciences, see, e.g., [24, 25].
For simplicity and ease of presentation, throughout the following we consider the case where the dynamics (1a) are subject to additive disturbances:
$f(x,u,w)=f_{\mathrm{a}}(x,u)+w$ (21)
with $w\in\mathcal{W}=\mathbb{R}^{n}$. Furthermore, we impose a Lipschitz condition on the nonlinear functions $f_{\mathrm{a}}$ and $h$ on $\mathcal{X}$.
Assumption 4
The functions $f_{\mathrm{a}}$ and $h$ are Lipschitz in $x\in\mathcal{X}$ uniformly for all $u\in\mathcal{U}$, i.e., there exist constants $L_{f},L_{h}>0$ such that
$|f_{\mathrm{a}}(x_{1},u)-f_{\mathrm{a}}(x_{2},u)|$ $\leq L_{f}|x_{1}-x_{2}|,$ (22)
$|h(x_{1},u)-h(x_{2},u)|$ $\leq L_{h}|x_{1}-x_{2}|$ (23)
for all $x_{1},x_{2}\in\mathcal{X}$ uniformly for all $u\in\mathcal{U}$.
Now, consider a batch of measured input-output data $d_{0:T}$ and the solution $z_{0:T}^{*}$ of the FIE problem $P(d_{0:T})$. To approximate $z_{0:T}^{*}$, we define the approximate estimator
$\hat{x}^{\mathrm{ae}}_{j}=\begin{cases}\zeta_{N}(j,d_{0:N}),&j\in\mathbb{I}_{[0:N/2]}\\ \zeta_{N}(N/2,d_{j-N/2:j+N/2}),&j\in\mathbb{I}_{[N/2+1:T-N/2-1]}\\ \zeta_{N}(N-T+j,d_{T-N:T}),&j\in\mathbb{I}_{[T-N/2:T]},\end{cases}$ (24)
where $\zeta_{N}$ is from (7). Note that $\hat{x}^{\mathrm{ae}}_{0:T}$ is a feasible state trajectory of system (1) under the dynamics (21); the corresponding disturbance inputs are
$\hat{w}_{j}^{\mathrm{ae}}=\hat{x}^{\mathrm{ae}}_{j+1}-f_{\mathrm{a}}(\hat{x}^{\mathrm{ae}}_{j},u_{j}),\ j\in\mathbb{I}_{[0,T-1]}.$ (25)
Let $\hat{z}^{\mathrm{ae}}_{j}:=(\hat{x}^{\mathrm{ae}}_{j},\hat{w}^{\mathrm{ae}}_{j})$, $j\in\mathbb{I}_{[0,T-1]}$, and $\hat{z}^{\mathrm{ae}}_{T}:=(\hat{x}^{\mathrm{ae}}_{T},0)$. The following lemma shows that our specific choice of $\hat{z}^{\mathrm{ae}}_{j}$ lies in a neighborhood of $z^{*}_{j}$ for all $j\in\mathbb{I}_{[0:T]}$.
Lemma 2
Suppose that Assumption 3 holds. Then, the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ satisfies $|\hat{z}^{\mathrm{ae}}_{j}-z^{*}_{j}|\leq 2\beta(C,N/2)$ for all $j\in\mathbb{I}_{[0:T]}$ and all possible data $d_{0:T}$.
Proof:
The sequence $\hat{z}^{\mathrm{ae}}_{0:T}$ as defined in (24) and (25) is essentially constructed from three parts, where for each part we can exploit the turnpike property from Assumption 3. In particular, for $j\in\mathbb{I}_{[0,N/2]}$, Assumption 3 with $\tau=0$ yields
$|\hat{z}^{\mathrm{ae}}_{j}-z^{*}_{j}|\leq\beta(C,N-j)\leq\beta(C,N/2),\ j\in\mathbb{I}_{[0,N/2]}.$
For $j\in\mathbb{I}_{[N/2+1,T-N/2-1]}$, we can use Assumption 3 with $\tau\in\mathbb{I}_{[1,T-N-1]}$:
$|\hat{z}^{\mathrm{ae}}_{j}-z^{*}_{j}|$ $\leq 2\beta(C,N/2),\ j\in\mathbb{I}_{[N/2+1,T-N/2-1]}.$
Finally, for $j\in\mathbb{I}_{[T-N/2,T]}$, we can use Assumption 3 with $\tau=T-N$:
$|\hat{z}^{\mathrm{ae}}_{j}-z^{*}_{j}|\leq\beta(C,N-T+j)\leq\beta(C,N/2),\ j\in\mathbb{I}_{[T-N/2,T]}.$
Combining these three cases yields the desired result. ∎
We are now able to state the main result of this section.
Theorem 3
Consider the system dynamics (21) and the special case of quadratic cost (4) and (5). Suppose that Assumptions 3 and 4 hold. Then, there exist ${\sigma}_{1},\sigma_{2}\in\mathcal{L}$ such that for any $\epsilon>0$, the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ satisfies the performance estimate
$J_{T}(\hat{x}^{\mathrm{ae}}_{0:T},\hat{w}^{\mathrm{ae}}_{0:T-1};d_{0:T})$
$\leq(1+\epsilon)V_{T}(d_{0:T})+\frac{1+\epsilon}{\epsilon}\left(T\sigma_{1}(N)+\sigma_{2}(N)\right)$ (26)
for any possible data $d_{0:T}$.
Before proving Theorem 3, we want to highlight some key properties of the performance estimate (26).
Remark 3 (Interpretation of the performance estimate)
  1. 1.
    The performance of the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ is approximately optimal on $\mathbb{I}_{[0:T]}$ with error terms that depend on the choices of $\epsilon$, $N$, and $T$.
  2. 2.
    In case of exponentially decaying sensitivity (i.e., Assumption 3 holds with $\beta(s,t)=ce^{-kt}$ for some $c,k>0$), which can be expected in many practical applications, the $\mathcal{L}$-functions $\sigma_{1}$ and $\sigma_{2}$ in (26) are of exponential form. Then, one can render the error terms arbitrarily small by suitable choices of $\epsilon$, $N$, and $T$. Furthermore, the performance $J_{T}$ converges exponentially to $V_{T}$ as $N$ increases. This behavior is also evident in the numerical example in Section V.
  3. 3.
    The performance estimate (26) grows linearly with $T$ and tends to infinity if $T$ approaches infinity. This property is to be expected (due to the fact that the turnpike is never exactly reached) and conceptually similar to (non-averaged) performance results in economic model predictive control, see, e.g., [17, Sec. 5], [10]. To make meaningful performance estimates in case $T\rightarrow\infty$, we analyze the averaged performance in Corollary 1 below.
  4. 4.
    Theorem 3 is stated for the practically relevant case of quadratic cost functions (4) and (5) for ease of presentation, but can easily be extended to more general cost functions that fulfill a weak triangle inequality.
Proof of Theorem 3: Using the definitions from (3)–(5), the performance criterion reads
$J_{T}(\hat{x}^{\mathrm{ae}}_{0:T},\hat{w}^{\mathrm{ae}}_{0:T-1};d_{0:T})=\sum_{j=0}^{T-1}l(\hat{x}^{\mathrm{ae}}_{j},\hat{w}^{\mathrm{ae}}_{j};d_{j})+g(\hat{x}^{\mathrm{ae}}_{T};d_{T})$
$=\sum_{j=0}^{T-1}|\hat{w}^{\mathrm{ae}}_{j}|_{Q}^{2}+|y_{j}-h(\hat{x}^{\mathrm{ae}}_{j},u_{j})|_{R}^{2}+|y_{T}-h(\hat{x}^{\mathrm{ae}}_{T},u_{T})|_{G}^{2}.$ (27)
From the definition of $\hat{x}^{\mathrm{ae}}_{j}$ in (24) and the fact that $f_{\mathrm{a}}(x_{j}^{*},u_{j})+w^{*}_{j}-x_{j+1}^{*}=0$ using (21), the square root of the input cost satisfies
$|\hat{w}^{\mathrm{ae}}_{j}|_{Q}$ $=|\hat{x}^{\mathrm{ae}}_{j+1}-x_{j+1}^{*}+f_{\mathrm{a}}(x_{j}^{*},u_{j})-f_{\mathrm{a}}(\hat{x}^{\mathrm{ae}}_{j},u_{j})+w^{*}_{j}|_{Q}$
$\leq|\hat{x}^{\mathrm{ae}}_{j+1}-x_{j+1}^{*}|_{Q}+L_{f}|x_{j}^{*}-\hat{x}^{\mathrm{ae}}_{j}|_{Q}+|w^{*}_{j}|_{Q}$ (28)
for all $j\in\mathbb{I}_{[0,T-1]}$, where in the last step we have used the triangle inequality and Assumption 4. Using Lemma 2, it follows that
$|\hat{x}^{\mathrm{ae}}_{j}-x^{*}_{j}|\leq|\hat{z}^{\mathrm{ae}}_{j}-z^{*}_{j}|\leq 2\beta(C,N/2)=:\sigma(N),\ j\in\mathbb{I}_{[0,T]},$
where we note that $\sigma\in\mathcal{L}$. Hence, from (28) we obtain
$|\hat{w}^{\mathrm{ae}}_{j}|_{Q}\leq|w_{j}^{*}|_{Q}+(1+L_{f})\lambda_{\max}(Q)\sigma(N),j\in\mathbb{I}_{[0,T-1]}.$
Squaring both sides and using the fact that for any $\epsilon>0$, it holds that $(a^{2}+b^{2})\leq(1+\epsilon)a^{2}+\frac{1+\epsilon}{\epsilon}b^{2}$ for all $a,b\geq 0$ then lets us conclude that
$|\hat{w}^{\mathrm{ae}}_{j}|^{2}_{Q}\leq(1+\epsilon)|w_{j}^{*}|^{2}_{Q}+\frac{1+\epsilon}{\epsilon}(1+L_{f})^{2}\lambda_{\max}(Q)^{2}\sigma(N)^{2}$ (29)
for each $j\in\mathbb{I}_{[0,T-1]}$. A similar reasoning for the fitting error (where we add $h(x_{j}^{*},u_{j})-h(x_{j}^{*},u_{j})=0$, $j\in\mathbb{I}_{[0,T]}$) yields
$|y_{j}-h(\hat{x}^{\mathrm{ae}}_{j},u_{j})|_{R}\leq|y_{j}-h(x_{j}^{*},u_{j})|_{R}+L_{h}\lambda_{\max}(R)\sigma(N)$
for all $j\in\mathbb{I}_{[0,T]}$. By squaring both sides and using the same argument that allowed us to obtain (29), we get
$|y_{j}-h(\hat{x}^{\mathrm{ae}}_{j},u_{j})|_{R}^{2}$ $\leq(1+\epsilon)|y_{j}-h(x_{j}^{*},u_{j})|_{R}^{2}$
$\quad+\frac{1+\epsilon}{\epsilon}L_{h}^{2}\lambda_{\max}(R)^{2}\sigma(N)^{2}$ (30)
for all $j\in\mathbb{I}_{[0:T]}$. A similar bound results for the terminal cost using the same arguments with $G$ instead of $R$:
$|y_{T}-h(\hat{x}^{\mathrm{ae}}_{T},u_{T})|_{G}^{2}$ $\leq(1+\epsilon)|y_{T}-h(x_{T}^{*},u_{T})|_{G}^{2}$
$\quad+\frac{1+\epsilon}{\epsilon}L_{h}^{2}\lambda_{\max}(G)^{2}\sigma(N)^{2}.$ (31)
The performance criterion (27) together with (29)–(31) and the definitions
$\sigma_{1}(s)$ $:=\big{(}(1{\,+\,}L_{f})^{2}\lambda_{\max}(Q)^{2}{+\,}L_{h}^{2}\lambda_{\max}(R)^{2}\big{)}\sigma(s)^{2},s\geq 0,$
$\sigma_{2}(s)$ $:=L_{h}^{2}\lambda_{\max}(G)^{2}\sigma(s)^{2},s\geq 0$
(with implies that $\sigma_{1},\sigma_{2}\in\mathcal{L}$) establishes the statement of this theorem and hence finishes this proof.  ∎
To establish an averaged performance estimate, we first need to show that $V_{T}$ grows at maximum linearly in $T$.
Lemma 3
Let $\mathcal{W}$ and $\mathcal{V}$ be compact. There exist $A,B>0$ such that $V_{T}(d_{0:T})\leq AT+B$ for any possible data $d_{0:T}$.
Proof:
Due to $\mathcal{W}$ and $\mathcal{V}$ being compact, there exist $C_{\mathrm{Q}},C_{\mathrm{R}},C_{\mathrm{G}}>0$ such that $|\hat{w}|^{2}_{Q}\leq C_{\mathrm{Q}}$, $|y-h(\hat{x},u)|_{R}^{2}\leq C_{\mathrm{R}}$, and $|y-h(\hat{x},u)|_{G}^{2}\leq C_{\mathrm{G}}$ for all $(\hat{x},\hat{w},u,y)\in\mathcal{X}\times\mathcal{W}\times\mathcal{U}\times\mathbb{R}^{p}$ such that $y-h(\hat{x},u)\in\mathcal{V}$. Hence, the claim holds with $A=C_{\mathrm{Q}}+C_{\mathrm{R}}$ and $B=C_{\mathrm{G}}$. ∎
Corollary 1 (Averaged performance)
Assume that $\mathcal{W}$ and $\mathcal{V}$ are compact. Then, the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ satisfies the averaged performance estimate
$\limsup\limits_{T\rightarrow\infty}\frac{1}{T}J_{T}(\hat{x}^{\mathrm{ae}}_{0:T},\hat{w}^{\mathrm{ae}}_{0:T-1};d_{0:T})$
$\leq\limsup\limits_{T\rightarrow\infty}\frac{1}{T}V_{T}(d_{0:T})+\epsilon A+\frac{1+\epsilon}{\epsilon}\sigma_{1}(N)$
for all possible data $d_{0:T}$, where $\epsilon>0$ and $\sigma_{1}\in\mathcal{L}$ are from Theorem 3 and $A>0$ is from Lemma 3.
Corollary 1 implies that the averaged performance of the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ is finite and approximately optimal with error terms that can be made arbitrarily small by suitable choices of $\epsilon$ and $N$.
IV-B Implications for moving horizon estimation
In online state estimation, the variable $T$ represents the forward (online) time that continuously grows (in the following, we use lowercase $t$ to indicate this case). Then, at each time instant $t\in\mathbb{I}_{\geq 0}$, one is interested in obtaining an accurate estimate of the current true unknown state $x_{t}$. Obviously, solving $P_{t}(d_{0:t})$ for the desired FIE solution $x^{*}_{t}$ is generally infeasible in practice since the problem size also continuously grows with time. Instead, MHE considers the truncated optimal estimation problem $P_{N}(d_{t-N:t})$ using the most recent data $d_{t-N:t}$ only, where the horizon length $N\in\mathbb{I}_{\geq 0}$ is fixed. More precisely, the MHE estimate at the current time $t\in\mathbb{I}_{\geq 0}$ is given by
$\hat{x}_{t}^{\mathrm{mhe}}=\begin{cases}\zeta_{t}(t,d_{0:t}),&t\in\mathbb{I}_{[0:N-1]}\\ \zeta_{N}(N,d_{t-N:t}),&t\in\mathbb{I}_{\geq N}.\end{cases}$ (32)
From our results in Section IV-A, we know that $\hat{x}_{t}^{\mathrm{mhe}}$ is actually the endpoint of some trajectory that is approximately optimal on $\mathbb{I}_{[0,t]}$ (namely, the sequence $\hat{x}^{\mathrm{ae}}_{0:t}$ from (24)), and hence indeed constitutes a meaningful estimate that is expected to be close to the desired (unknown) FIE solution ${x}_{t}^{*}=\zeta_{t}(t,d_{0:t})$ if $N$ is sufficiently large. In particular, under the turnpike property from Assumption 3, we can explicitly bound their difference by
$|\hat{x}_{t}^{\mathrm{mhe}}-x^{*}_{t}|=|\zeta_{N}(N,d_{t-N:t})-\zeta_{t}(t,d_{0:t})|\leq\beta(C,N)$
for all $t\in\mathbb{I}_{\geq N}$, which in fact can be made arbitrarily small by a suitable choice of $N$ (note that for $t\in\mathbb{I}_{[0,N-1]}$, the FIE and MHE solutions coincide anyway).
However, we also want to emphasize that the MHE sequence $\hat{x}^{\mathrm{mhe}}_{0:t}$ (32) is not approximately optimal on the whole interval $\mathbb{I}_{[0,t]}$ (and therefore turns out to be rather unsuitable for offline estimation), because it is the concatenation of solutions of truncated problems that lie on the leaving arc for all $j\in\mathbb{I}_{[0,t-1]}$, which might actually be far away from the turnpike (the optimal FIE solution). This effect is also evident in the simulation example in the following section.
V Numerical example
We consider the system
$x_{1}^{+}$ $=x_{1}+t_{\Delta}(-2k_{1}x_{1}^{2}+2k_{2}x_{2})+u_{1}+w_{1},$
$x_{2}^{+}$ $=x_{2}+t_{\Delta}(k_{1}x_{1}^{2}-k_{2}x_{2})+u_{2}+w_{2},$
$y$ $=x_{1}+x_{2}+v,$
[…]
In practice, one is usually also interested in the accuracy of the estimated state trajectory with respect to the real unknown system trajectory $x_{0:T}$. Therefore, we compare the sum of the normed errors222 For a given sequence $\hat{x}_{0:T}$, we define $\mathrm{SNE}(\hat{x}_{0:T}):=\sum_{j=0}^{T}|\hat{x}_{j}-x_{j}|$. (SNE) of the FIE solution $x^{*}_{0:T}$, the approximate estimator $\hat{x}^{\mathrm{ae}}_{0:T}$, and MHE $x^{\mathrm{mhe}}_{0:T}$ for different sizes $N$ of the truncated problems. The corresponding results in Table I show qualitatively the same behavior as in the previous performance analysis. In particular, the FIE solution yields the most accurate estimates with the lowest SNE. The proposed approximate estimator yields much higher SNE for small horizons (SNE increase of 54 % for $N=40$ compared to FIE), but improves very fast as $N$ increases, and exponentially converges to the SNE of FIE. On the other hand, the SNE of MHE improves much slower, and is particularly much worse than that of the FIE solution and the proposed approximate estimator (for $N\geq 70$).
Figure 2: Performance $J_{T}$ of the approximate estimator $\hat{z}^{\mathrm{ae}}_{0:T}$ (blue) and MHE $\hat{z}^{\mathrm{mhe}}_{0:T}$ (green) for different lengths $N$ of the truncated problems $P_{N}$ compared to the optimal performance of the FIE solution $z^{*}_{0:T}$ (red).
TABLE I: SNE for the proposed approximate estimator and MHE.
  Problem length $N$ Approximate estimator MHE
   40 92.603  (+52.7 %) 71.334  (+17.6 %)
   70 63.417  (+4.6 %) 66.588  (+9.8 %)
  100 61.387  (+1.2 %) 65.429  (+7.9 %)
  130 61.025  (+0.6 %) 64.771  (+6.8 %)
  160 60.805  (+0.3 %) 65.007  (+7.2 %)
  • Values in parentheses indicate the relative increase in the SNE compared
  • to the optimal FIE solution $x^{*}_{0:T}$ (which achieves SNE = 60.638).
VI Conclusion
We have shown that the optimal FIE solution serves as turnpike for the solutions of truncated state estimation problems involving a subset of the measurement data only. We proposed an estimator for approximating the optimal FIE solution using a sequence of truncated problems (which can be solved in parallel), and we showed that the resulting performance is approximately optimal. The numerical example illustrates that the performance of our proposed approximate estimator exponentially converges to the optimal performance if the problem length is increased (which is not the case for MHE). In future work, we will investigate how these results can be used to improve the performance of MHE.